Let the harmonic mean and geometric mean of two positive  numbers be the ratio 4 : 5.
Then the two number are in the ratio=

Grade:9

1 Answers

Arun Kumar IIT Delhi
10 years ago
Hello Student,
\\\sqrt{ab}={2ab \over a+b}*{5 \over 4} \\=>4(a+b)=10\sqrt{ab} \\=>4(\sqrt{a \over b}+\sqrt{b \over a})=10 \\=>t+1/t=5/2
Please solve for t

Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty